8. Solve the following initial value problem: y - 6y + 13y = 0, y(a) = b, y(c) = d (Note: a, b,c,d will be an a=1 b=7 c=8
Answer #1

ч – 68 +1 234- о. (4) - Ы. ч) = 4 (1) . )- →y у — 6s - Ina - ҳ Алді «%я Equaton - - Gм + 1 = 6+ \з, -52 3 Ә 6 + 4) м = 4 +3 colib-2Bint) + 3 Sinib+ 2 C0216) = ge 24 [ora at sinac ca 7. e 3 >(138in1b.coba +200816.6082_360816.8ina + 2 sinlb.sina] =- > (ß cosio. Sin 2 -2 sinit. Sina-3 finit. Cosa-2 Casio Coss] ge 24 eina - 7e? (38anitta cosib) <,..(-3 sinia - 2 cosial gē3